OpenStudy (anonymous):

10^6 in standard form? fan and medal

OpenStudy (skullpatrol):

Define "standard" form please?

OpenStudy (anonymous):

10000000

OpenStudy (skullpatrol):

What do you mean?

OpenStudy (anonymous):

or 10 with 6 zeros

OpenStudy (skullpatrol):

Is it the same as scientific notation?

OpenStudy (anonymous):

like put it in form i dont know if its 1 with six 0's or 10 with 6 more 0's

OpenStudy (anonymous):

10 with 6 more

OpenStudy (anonymous):

1000000 is the answer if you were wanting to use 10^6 as standard form

OpenStudy (anonymous):

thank you menakeeta

OpenStudy (skullpatrol):

$$10^6 = 1.0*10^6$$

OpenStudy (anonymous):

menakeeta already did it but thank you all.

OpenStudy (anonymous):

no because 10^6 is a 1 with 6 0's

OpenStudy (anonymous):

yeah thats what i got

OpenStudy (anonymous):

your welcome(:

OpenStudy (anonymous):

do it on a calculator

OpenStudy (skullpatrol):

$$10^6=1.0*10^6=1,000,000$$

OpenStudy (anonymous):

thanks guys have a nice day.